# Energy-Efficient Balanced Flow Control Achieved through Optimization of Synthetic Jet Placement Based on Deep Reinforcement Learning

### v1

@@ 4.2.1 Scenario B: Control Performance of Multiple Jets Control
Figure 17: Evaluation of flow control performance for a square cylinder using single and multiple synthetic jets configurations. (a) Reward curves during the training process for different jet placement configurations. (b) Drag coefficient ($C_{D}$) as a function of non-dimensional time for baseline and controlled flows. (c) Lift coefficient ($C_{L}$) as a function of non-dimensional time for baseline and controlled flows.
We further designed experiments to implement flow control using multiple pairs of synthetic jets simultaneously. The results of DRL training and its impact on flow control for a square cylinder with various jet configurations are illustrated in figure 17. For comparison, results using a single pair of synthetic jets are also provided. The evolution of the reward functions during training for different jet configurations is shown in figure 17(a). The reward function serves as a critical metric for evaluating the effectiveness of the DRL agent in optimizing flow control strategies. In this figure, $J^{1}_{2}J^{2}_{2}$, $J^{1}_{3}J^{2}_{3}$, and $J^{1}_{4}J^{2}_{4}$ represent specific jet positions on the square cylinder.
Across all four configurations, the reward functions increase as training progresses, indicating that the DRL agent is successfully learning and refining the flow control strategies. Initially, the reward values are low, reflecting the early stages of learning when the agent explores various control strategies. As the number of episodes increases, the reward function for each jet configuration shows steady improvement, particularly between 500 and 2000 episodes. The gradual rise in the reward values suggests that the DRL algorithm is progressively enhancing its strategy, leading to improved control performance. Notably, the reward function for the $J^{1}_{3}J^{2}_{3}$ configuration exhibits faster growth, indicating that this setup is particularly effective in achieving the control objectives. The quicker convergence suggests that the agent is learning the optimal control strategy more efficiently with this configuration. Conversely, the more gradual increase in reward functions observed when using multiple pairs of synthetic jets suggests a more incremental learning process. The slower convergence may result from increased complexity in the fluid dynamics or suboptimal jet placement for these specific control tasks.
@@
Over non-dimensional time for various jet configurations, compared to a baseline scenario without active control, the lift coefficient is illustrated in figure 17(c). The baseline $C_{L}$ without control exhibits periodic oscillations. The activation of synthetic jets significantly reduces both the amplitude and frequency of these oscillations, stabilizing the lift coefficient closer to zero or the desired level. This suppression of lift fluctuations helps create a more stable flow field by mitigating the unsteady forces acting on the cylinder. Specifically, the configuration with jets positioned at $J_{3}^{1},J_{3}^{2}$ shows a more pronounced reduction in lift compared to those at $J_{2}^{1},J_{2}^{2}$ or $J_{4}^{1},J_{4}^{2}$. When two jet pairs ($J_{2}^{1},J_{2}^{2}$ and $J_{3}^{1},J_{3}^{2}$) are activated simultaneously, the reduction in lift is similar to that achieved with $J_{3}^{1},J_{3}^{2}$ alone, indicating that the additional jet pair does not significantly enhance control. When all three pairs of jets ($J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$) are activated simultaneously, the oscillations in the lift coefficient become more pronounced, indicating increased flow instability. While activating all three jet pairs enhances drag reduction, it also introduces additional instability in the lift coefficient. This analysis demonstrates that the $J_{3}^{1},J_{3}^{2}$ configuration provides the most efficient and effective flow control, as evidenced by rapid reward function convergence, significant drag reduction, and stabilized lift. Although the simultaneous activation of two jet pairs ($J_{2}^{1},J_{2}^{2}$ and $J_{3}^{1},J_{3}^{2}$) yields similar control performance, it does not achieve the same level of convergence efficiency as using $J_{3}^{1},J_{3}^{2}$ alone. The activation of all three jet pairs, while improving drag reduction, leads to increased flow instability, which will be further analyzed in the next section in relation to the energy consumption required for flow control.
@@ 4.2.2 Scenario B: Energy Consumption with Simultaneous Control (Two Pairs and Three Pairs )
The external energy consumption over non-dimensional time for flow control with two (a) and three (b) synthetic jets pairs around a square cylinder is depicted in figure 18. When comparing the single-pair synthetic jets control in figure 15 to the multi-pair configuration in figure 18, it is evident that the energy consumption in the multi-jet scenario is reduced by an order of magnitude relative to the single-jet case. Where two pairs of synthetic jets, $J_{2}^{1},J_{2}^{2}$ and $J_{3}^{1},J_{3}^{2}$, are simultaneously employed, significant peaks in action values for both configurations are shown in the initial phase in figure 18(a), indicating a high energy demand at the onset of flow control. This peak reflects the system response to the abrupt introduction of control forces generated by the jets. As time progresses, the action values stabilize, signifying that the energy required to maintain control has reached a steady state. Notably, the energy demand for the jets positioned at $J_{3}^{1},J_{3}^{2}$ consistently remains lower than that for $J_{2}^{1},J_{2}^{2}$, highlighting the greater energy efficiency of the $J_{3}$ configuration. The sustained lower action values suggest that this configuration achieves the desired flow control with less external energy input, making it a more efficient option for long-term operation.
The control performance is illustrated in figure 18(b) when three pairs of jets ($J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$) are employed simultaneously. The action values indicate the energy efficiency of various jet configurations in controlling the flow around the square cylinder, with lower sustained values signifying more energy-efficient setups. Similar to the two-jet configuration, the three-jet setup also exhibits peak action values during the initial phase of control, indicative of the high initial energy demand required to establish effective control. As the system stabilizes, the action values across all jet pairs settle into a steady state. Among the three-jet configurations, the jets positioned at $J_{4}^{1},J_{4}^{2}$ demonstrate the highest sustained action values, indicating a higher energy requirement to maintain effective flow control. In contrast, the jets at $J_{3}^{1},J_{3}^{2}$ consistently show lower energy consumption compared to those at $J_{4}^{1},J_{4}^{2}$ and J21,J2

### v5

@@ 4.2.1 Scenario B: Control Performance of Multiple Jets Control
Figure 16: Evaluation of flow control performance for a square cylinder using single and multiple synthetic jets configurations. (a) Reward curves during the training process. (b) $C_{D}$ over non-dimensional time. (c) $C_{L}$ over non-dimensional time.
To further explore the influence of actuator placement on flow control performance, additional experiments were conducted by simultaneously employing multiple pairs of synthetic jets. The corresponding results, including the training process of the DRL agent and its flow control effectiveness for a square cylinder with various jet configurations, are presented in figure 16. For comparison, the performance achieved using a single pair of synthetic jets is also included. Figure 16(a) illustrates the temporal evolution of the reward function throughout the training process for different jet configurations. In particular, the configurations $J^{1}_{2}J^{2}_{2}$, $J^{1}_{3}J^{2}_{3}$, and $J^{1}_{4}J^{2}_{4}$ denote specific synthetic jet locations on the square cylinder surface, and their comparative performance underscores the impact of jet placement on learning efficiency and flow control capability.
Across all four configurations, the reward functions increase as training progresses, indicating that the DRL agent is successfully learning and refining the flow control strategies. Initially, the reward values are low, reflecting the early stages of learning when the agent explores various control strategies. As the number of episodes increases, the reward function for each jet configuration shows steady improvement, particularly between 500 and 2000 episodes. The gradual rise in the reward values suggests that the DRL algorithm is progressively enhancing its strategy, leading to improved control performance. Crucially, the reward function for the $J^{1}_{3}J^{2}_{3}$ configuration exhibits faster growth, indicating that this setup is particularly effective in achieving the control objectives. The quicker convergence suggests that the agent is learning the optimal control strategy more efficiently with this configuration. Conversely, the more gradual increase in reward functions observed when using multiple pairs of synthetic jets suggests a more incremental learning process. The slower convergence may result from increased complexity in the fluid dynamics or suboptimal jet placement for these specific control tasks.
@@
Over non-dimensional time for various jet configurations, compared to a baseline scenario without active control, the lift coefficient is illustrated in figure 16(c). The baseline $C_{L}$ without control exhibits periodic oscillations. The activation of synthetic jets significantly reduces both the amplitude and frequency of these oscillations, stabilizing the lift coefficient closer to zero or the desired level. This suppression of lift fluctuations helps create a more stable flow field by mitigating the unsteady forces acting on the cylinder. Specifically, the configuration with jets positioned at $J_{3}^{1},J_{3}^{2}$ shows a more pronounced reduction in lift compared to those at $J_{2}^{1},J_{2}^{2}$ or $J_{4}^{1},J_{4}^{2}$. When two jet pairs ($J_{2}^{1},J_{2}^{2}$ and $J_{3}^{1},J_{3}^{2}$) are activated simultaneously, the reduction in lift is similar to that achieved with $J_{3}^{1},J_{3}^{2}$ alone, indicating that the additional jet pair does not significantly enhance control.
When all three pairs of synthetic jets ($J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$) are activated simultaneously, the lift coefficient exhibits intensified oscillations, signifying increased flow instability. Although this configuration leads to notable drag reduction, the resulting unsteady aerodynamic forces may undermine the overall robustness of the control strategy. This highlights a fundamental trade-off between achieving lower drag and maintaining stable lift forces. A comparative analysis of different jet configurations further elucidates this trade-off. Among the tested configurations, the $J_{3}^{1},J_{3}^{2}$ placement proves to be the most effective, demonstrating rapid reward function convergence, significant drag reduction, and stabilized lift dynamics. In contrast, while the simultaneous activation of two jet pairs ($J_{2}^{1},J_{2}^{2}$ and $J_{3}^{1},J_{3}^{2}$) yields a comparable level of drag mitigation, it fails to match the convergence efficiency achieved by the $J_{3}^{1},J_{3}^{2}$ configuration alone. Furthermore, while the deployment of all three jet pairs amplifies drag reduction, the induced lift fluctuations underscore the trade-off between aerodynamic performance and flow stability. This interplay between control effectiveness and induced instabilities will be further examined in the subsequent section, with a specific focus on the energy expenditure associated with different jet actuation strategies.
@@ 4.2.2 Scenario B: Energy Consumption with Simultaneous Control (Two Pairs and Three Pairs)
The external energy consumption over non-dimensional time for flow control with two (a) and three (b) synthetic jets pairs around a square cylinder is depicted in figure 17. When comparing the single-pair synthetic jets control in figure 14 to the multi-pair configuration in figure 17, it is evident that the energy consumption in the multi-jet scenario is reduced by an order of magnitude relative to the single-jet case. Where two pairs of synthetic jets, $J_{2}^{1},J_{2}^{2}$ and $J_{3}^{1},J_{3}^{2}$, are simultaneously employed, significant peaks in action values for both configurations are shown in the initial phase in figure 17(a), indicating a high energy demand at the onset of flow control. This peak reflects the system response to the abrupt introduction of control forces generated by the jets. As time progresses, the action values stabilize, signifying that the energy required to maintain control has reached a steady state.
The total energy consumption of the synthetic jets located at $J_{3}^{1}$ and $J_{3}^{2}$ is significantly lower than that of the jets positioned at $J_{2}^{1}$ and $J_{2}^{2}$, highlighting the superior energy efficiency of the $J_{3}$ configuration. This indicates that the $J_{3}^{1}$ and $J_{3}^{2}$ jet configuration achieves the desired flow control with reduced external energy input, making it a more effective option for long-term operation.
Figure 17(b) presents the control performance when three pairs of synthetic jets ($J_{2}^{1},J_{2}^{2}$; $J_{3}^{1},J_{3}^{2}$; and $J_{4}^{1},J_{4}^{2}$) are simultaneously deployed around the square cylinder. The control action quantified by the time integral of actuation signals, reflects the total energy expenditure required to implement flow control via synthetic jets. A lower cumulative actuation input corresponds to reduced external energy consumption, thereby indicating enhanced energy efficiency. As in the two-jet scenarios, all jets exhibit elevated actuation levels during the initial transient phase, reflecting the higher energy demand required to rapidly suppress vortex shedding and stabilize the wake. Once the control effect is established, the actuation levels approach a quasi-steady state.
Among the three configurations, the jets located at $J_{4}$ maintain the highest sustained action values, indicating a greater continuous energy requirement for maintaining effective control. In contrast, the jets at $J_{3}$ consistently exhibit lower energy expenditure throughout the control horizon. This observation not only corroborates their superior performance in the two-jet configuration but also highlights their robustness in more complex control setups. The persistent energy efficiency of the $J_{3}$ configuration across both two- and three-jet arrangements suggests that this location—likely corresponding to critical regions such as separation or reattachment points—offers an optimal balance between control authority and energy cost. These findings are consistent with prior results obtained from single-jet analysis, where jets at $J_{3}$ demonstrated effective suppression of unsteady wake dynamics with minimal actuation. Overall, the analysis emphasizes the critical role of jet placement in achieving energy-efficient flow control, with $J_{3}$ emerging as the most favorable location for long-term, low-energy operation.
5 Discussion: the physical insight of flow control strategy
5.1 Flow separation phenomenon
Section 3 and section 4 provide a detailed examination of the effects of various jet configurations on flow control performance. To further analyze the underlying mechanisms of flow control in relation to fluid dynamics, we present a comprehensive discussion of the control performance for both circular and square cylinder flows. When $Re=100$, the flow around the circular cylinder exhibits a stable laminar characteristic with well-defined separation points, as shown in figure 18. Flow separation occurs symmetrically on both sides of the cylinder, approximately 105° from the forward stagnation point. At this location, the adverse pressure gradient causes the boundary layer to detach from the surface, marking the onset of flow separation. This detachment initiates the formation of a wake behind the cylinder, where alternating vortices periodically shed, ultimately leading to the development of the characteristic von Kármán vortex street downstream.
Figure 18: Streamline pattern for the baseline flow at $Re=100$. (a) circular cylinder and (b) square cylinder.
In contrast, for a square cylinder at the same Reynolds number, the sharp changes in geometry result in flow separation occurring at the four sharp corners of the body. The boundary layer detaches immediately at these edges, forming distinct vortices. Especially, large separation bubbles form near the rear corners, creating a pronounced separation zone in the wake. The differing nature of flow separation around circular and square cylinders necessitates distinct control strategies, posing unique challenges when applying reinforcement learning algorithms for flow control.
5.2 Flow control mechanism
Through comparative analysis, the optimal position for synthetic jets around a circular cylinder is found to be at a 105° angle from the front stagnation point. At this position, drag and lift are significantly reduced, vortex shedding is fully suppressed, and energy consumption is minimized. This result aligns with the findings of Wang et al. (2024) at $Re=100$, which demonstrated superior control performance using a sensor at 105°. The results further confirm that the 105° configuration offers the best balance between energy efficiency and flow stability, as detailed in section 3.
Figure 19: (a) Velocity contours of flow around a cylinder in the baseline case. (b) Velocity contours of the controlled flow with synthetic jets positioned at a 105° azimuthal angle. (c) Streamlines around the cylinder in the baseline case. (d) Streamlines around the cylinder with active synthetic jets at a 105° azimuthal angle. (e) Streamlines around the cylinder after flow control is completed with synthetic jets activated at a 105° azimuthal angle.
Figure 19 illustrates the effectiveness of flow control achieved by synthetic jets positioned at 105° ($J_{2}^{1}$ and $J_{2}^{2}$), presenting instantaneous streamline patterns at three representative time instances: prior to, during, and after the activation of control. In the baseline case shown in figure 19(a), a classical Kármán vortex street is observed downstream of the cylinder, characteristic of unsteady wake dynamics. Upon activation of the synthetic jets, as depicted in figure 19(b), the recirculation region on the leeward side of the cylinder expands significantly, and vortex shedding is completely suppressed. Furthermore, figure 19(c) reveals the baseline flow’s separation zone and the development of a large, closed vortex structure on the rear side of the cylinder, which is effectively mitigated through synthetic jet control.
When the jet positioned at the 105° azimuth is activated, it operates through alternating ejection and suction of the surrounding fluid, effectively altering the local flow without introducing any new mass into the system, as shown in figure 19 (d). The synthetic jets located at $J_{2}^{1}$ and $J_{2}^{2}$ consistently satisfy the condition $Q_{1}+Q_{2}=0$, ensuring that the net mass flow in the system remains zero. However, the momentum and hydrodynamic impulse generated by the jets are not zero. The interaction between the synthetic jets and the crossflow leads to the formation of new vortex pairs. These high-energy vortices convect over the cylinder, locally modifying the streamlines around its surface, as shown in figure 19 (e). As a result, the apparent aerodynamic shape of the cylinder is significantly improved, along with its aerodynamic characteristics. By altering the local streamlines and inducing noticeable surface modifications, synthetic jets effectively control the flow, reducing drag and improving overall aerodynamic efficiency. Smith and Glezer (1998) and Amitay et al. (1997) experimentally demonstrated the formation, evolution, and control mechanisms of synthetic jets in flow control. The use of a DRL-based approach reveals a dynamically evolving flow control process that aligns with their findings. This consistency further validates the effectiveness of synthetic jets in altering flow behavior and enhancing aerodynamic performance through active flow control strategies.
In section 4, the training results for various jet configurations around the square cylinder demonstrate that positioning synthetic jets near the rear corners of the cylinder yields the optimal control performance, achieving both low energy consumption and high flow control efficiency. In this configuration, drag reduction and lift suppression are significantly enhanced, while vortex shedding in the wake is entirely suppressed with minimal external energy input. Further analysis of the separation points for both square and circular cylinders, as shown in section 18, combined with the results from section 3 and section 4, indicates that placing synthetic jets near the flow separation points is critical for effective flow control. Only when the jets are positioned in close proximity to these separation points can vortex shedding be fully suppressed.
Figure 20: (a) Velocity contours of flow around a square cylinder in the baseline case. (b) Streamlines around the square cylinder in the baseline. (c) Velocity contours of the controlled flow with jets positioned at the $J_{3}$ location. (d) Streamlines around the square cylinder after flow control with jets positioned at the $J_{3}$ location.
The comparison between the baseline flow and the controlled flow, with jets positioned at the rear corners ($J_{3}^{1}$ and $J_{3}^{2}$), is illustrated in figure 20. The placement of the synthetic jets near the rear corners effectively suppresses vortex formation and stabilizes the wake flow. Streamline patterns around the square cylinder reveal significant local alterations near the rear corners due to the action of the synthetic jets. In the controlled flow, the flow separation regions on both sides of the square cylinder become more defined, and two stable, closed separation bubbles form on the cylinder’s rear face. The interaction between synthetic jets and the external flow induces significant changes in local flow patterns and streamlines, facilitating linear momentum transfer without mass addition. The alternating phases of blowing and suction generate vortices that enhance boundary layer mixing, supplying the necessary energy to overcome adverse pressure gradients and enabling complete flow reattachment around the square cylinder.
6 CONCLUSIONS
We use DRL algorithms to train synthetic jet-based flow control strategies for circular and square cylinders. The agent learns to optimize jet actuation to minimize drag and suppress instabilities while balancing energy efficiency and control effectiveness. By systematically evaluating single- and multi-jet configurations, we identify optimal jet placements and provide qualitative guidelines based on hydrodynamic analyses.
  1. 1.
    Flow around a circular cylinder scenario. The single-jet training results indicate that positions $J_{1}$ and $J_{5}$ show slower and lower reward function convergence, suggesting lower training effectiveness and energy efficiency. In contrast, positions $J_{2}$, $J_{3}$, and $J_{4}$ achieve rapid convergence and superior control strategy acquisition. Despite all positions achieving an 8% drag reduction and 99% lift coefficient suppression, $J_{1}$ and $J_{5}$ require three times the energy of $J_{2}$ to reach the same performance level. For the control strategy involving the simultaneous activation of two pairs of synthetic jets, the configuration utilizing jets at positions $J_{2}$ and $J_{4}$ demonstrates higher energy efficiency compared to the configuration using jets at position $J_{3}$. When three pairs of synthetic jets are activated simultaneously, the control strategy with jets located at $J_{2}$ demonstrates the highest energy efficiency, compared to those at $J_{4}$ or $J_{3}$.
  2. 2.
    Flow around a square cylinder scenario. When controlling a single pair of synthetic jets, placement at $J_{3}$ yields the best performance, achieving 14.4% drag reduction with only 1.7% of the inlet flow rate—demonstrating both high control effectiveness and energy efficiency. In contrast, jets at $J_{1}$ require 44% of the flow rate for just 12.1% drag reduction, resulting in higher energy cost and lower stability. Using dual jet pairs at $J_{2}$ and $J_{3}$ improves force stabilization and achieves performance comparable to $J_{3}$ alone, but with slower reward convergence. When all three jet pairs $J_{2}$, $J_{3}$, $J_{4}$ are active, $J_{3}$ still dominates in efficiency and control effectiveness, highlighting its optimal placement.
  3. 3.
    Mechanism of energy-efficient flow control. DRL training results reveal that placing synthetic jets near 105° on a circular cylinder and at the rear corners of a square cylinder yields optimal energy efficiency and flow control performance. Positioning jets at separation points enhances momentum exchange and induces beneficial vortex structures, promoting boundary layer reattachment without mass addition. This strategic alignment with key flow dynamics significantly improves control effectiveness, stability, and energy efficiency.
The effectiveness of DRL in flow control lies in its ability to autonomously learn optimal strategies through interaction with complex, nonlinear environments. By capturing intricate state–action relationships, DRL enables exploration of control solutions beyond the scope of conventional methods. This study demonstrates that strategic placement of synthetic jets allows for a favorable trade-off between control performance and energy efficiency. Coordinated jet configurations yield synergistic effects, enhancing overall control efficacy. The findings underscore the pivotal role of jet placement in achieving robust, energy-efficient flow control.
7 APPENDIX I
To ensure that the numerical results are independent of the mesh resolution, a mesh independence study is conducted for both circular and square bluff bodies. The computational domain is discretized using three levels of mesh density: coarse, medium, and fine. The total number of grid cells for each case is reported in table 2. The drag coefficient $C_{D}$, lift coefficient $C_{L}$, and Strouhal number $St$ are selected as key parameters for evaluation.
Table 2: Mesh independence test for the flow around a bluff body at $Re=100$.
Bluff Configuration Mesh $C_{D,mean}$ $C_{D,max}$ $C_{L,max}$ $St$
Coarse 10,540 3.224 3.242 1.052 0.304
cylinder Medium 18,484 3.205 3.225 0.990 0.300
Fine 26,500 3.207 3.228 0.992 0.301
Schäfer et al. (1996) - - 3.220–3.240 0.990–1.010 0.295–0.305
Coarse 8465 1.532 1.562 0.326 0.145
square Medium 23 264 1.548 1.559 0.321 0.142
Fine 35 179 1.549 1.561 0.320 0.141
Singh et al. (2009) - 1.510 - - 0.147
Sen et al. (2011) - 1.529 - - 0.145
For the circular cylinder, the mean drag coefficient, maximum drag coefficient, and maximum lift coefficient all exhibit small variations and converge towards consistent values with increasing mesh resolution. The Strouhal number also remains stable across different mesh resolutions. These results are in close agreement with the reference benchmark study by Schäfer et al., further validating the numerical setup. For the square cylinder, a similar trend of convergence is observed for the mean drag coefficient and Strouhal number, with minimal changes across different mesh resolutions. The final values obtained are in excellent agreement with prior studies, such as Singh et al. and Sen et al.. Based on these observations, the medium mesh resolution is deemed sufficient for achieving accurate and computationally efficient results. Therefore, all subsequent simulations are conducted using the medium mesh to balance computational cost and solution accuracy.
8 APPENDIX II
The performance of two policy-based reinforcement learning algorithms in active flow control around a circular cylinder is comprehensively compared in figure 21. As shown in figure 21(a), the control strategy trained using the PPO algorithm consistently achieves the highest cumulative reward throughout the training process, exhibiting faster convergence and reduced fluctuations. This indicates superior learning efficiency and control performance. In contrast, the three variants of the vanilla policy gradient algorithm (curves a, b, and d) converge more slowly and exhibit significantly higher variance, reflecting unstable learning dynamics and suboptimal control effectiveness.
Figure 21: Performance evaluation of different jets configurations for flow control around a cylinder. (a) Reward curves over training episodes for various jets positions. (b) $C_{D}$ over non-dimensional time, comparing the baseline scenario with different jets configurations. (c) $C_{L}$ over non-dimensional time, comparing the baseline scenario with different jets.
The time-resolved aerodynamic performance of the learned controllers is further assessed in terms of the drag and lift coefficients. Figure 21(b) shows the evolution of the drag coefficient $C_{D}$, where all control strategies yield a gradual decrease in $C_{D}$ toward a lower steady-state value compared to the uncontrolled baseline. The PPO-based controller achieves the most rapid and pronounced drag reduction, with $C_{D}$ dropping below 2.98. This result highlights the efficacy of PPO in mitigating wake dynamics and suppressing vortex-induced drag. Figure 21(c) presents the temporal evolution of the lift coefficient $C_{L}$. The PPO-based strategy exhibits the fastest decay of $C_{L}$ oscillations, achieving near-zero mean fluctuations within a short time window, thereby indicating improved wake stabilization. In contrast, the Vanilla PG-based controllers show slower decay rates and larger residual oscillations, suggesting inferior suppression of unsteady lift forces.
In summary, the PPO algorithm demonstrates clear advantages over the Vanilla PG approaches in both training stability and flow control performance. It facilitates faster policy acquisition, more effective drag suppression, and enhanced lift stabilization. The consistent improvements observed in reward convergence, drag reduction, and lift damping collectively confirm the effectiveness and robustness of PPO-based strategies for active flow control applications.
9 APPENDIX III
To enhance the reproducibility of the present study, the principal parameters employed in both the CFD simulations and the DRL training procedure are summarized in table 3. The simulations are conducted at a Reynolds number of $Re=100$, with a non-dimensional time step of $5\times 10^{-4}$. To ensure efficient data generation and throughput, 60 parallel CFD environments are instantiated, each bound to a single CPU core, leading to a total allocation of 61 CPU cores. The deep reinforcement learning agent, based on the PPO algorithm, is trained over 3000 episodes. The relevant parameters of the PPO algorithm are summarized in the table.
Table 3: Parameters of both the CFD simulation and the DRL algorithm.
Parameter Symbol Value
Reynolds number $Re$ 100
Numerical time step (non-dimensional) dt $5\times 10^{-4}$
Parallel CFD environments - 60
CPUs/environment - 1
Total CPUs - 61
Episode number - 3000
Learning rate lr 0.001
Discount factor $\gamma$ 0.97
Policy network $\pi_{\theta}$ 512*512
Policy Ratio Clipping $\epsilon$ 0.3
Optimizer - Adam
Batch size - 64